Shift report 20030318 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Mar 18 2003 - 00:05:31 EST

  • Next message: Apache: "Shift report 20030318 00:00-08:00"
    ------------------------------------------------------------
    Date: Tue Mar 18 00:05:30 2003, Shift: 16:00-24:00
    Supervisor: CEJ, with: 
    Log book#14 pages 67 - 67
    
    Shift Summary:
    ------------------------------------------------------------
    1) No beam during shift. Several succesful injection of
       deuterons, but the gold just wouldn't stay in the ring.
    
    2) Only ped, TPC ped and DVM runs taken.
    
    3) Problems with the water pumps to the magnets, but they
       have now been stable since repair around 21:00. No 
       other problems.
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Tue Mar 18 2003 - 00:06:54 EST